 1 \ifx\LdfInit\@undefined
 2 \def\LdfInit{%
 3  \chardef\atcatcode=\catcode`\@
 4  \catcode`\@=11\relax
 5  \input babel.def\relax
 6  \catcode`\@=\atcatcode \let\atcatcode\relax
 7  \LdfInit}
 8 \fi
 9 ⟨+american⟩\def\CurrentOption{american}
10 ⟨+british⟩\def\CurrentOption{british}
11 ⟨+english⟩\def\CurrentOption{english}

12 ⟨american | british | english⟩\input english.ldf\relax

13 ⟨+bahasa⟩\input bahasa.ldf\relax
14 ⟨+breton⟩\input breton.ldf\relax
15 ⟨+catalan⟩\input catalan.ldf\relax
16 ⟨+croatian⟩\input croatian.ldf\relax
17 ⟨+czech⟩\input czech.ldf\relax
18 ⟨+danish⟩\input danish.ldf\relax
19 ⟨+afrikaans⟩\def\CurrentOption{afrikaans}
20 ⟨+dutch⟩\def\CurrentOption{dutch}
21 ⟨+afrikaans | dutch⟩\input dutch.ldf\relax
22 ⟨+esperant⟩\input esperant.ldf\relax
23 ⟨+estonian⟩\input estonian.ldf\relax
24 ⟨+finnish⟩\input finnish.ldf\relax
25 ⟨+francais⟩\def\CurrentOption{francais}
26 ⟨+french⟩\def\CurrentOption{french}
27 ⟨+francais | french⟩\input frenchb.ldf\relax
28 ⟨+galician⟩\input galician.ldf\relax
29 ⟨+austrian⟩\def\CurrentOption{austrian}
30 ⟨+german⟩\def\CurrentOption{german}
31 ⟨+germanb⟩\def\CurrentOption{german}

32 ⟨+austrian | german | germanb⟩\input germanb.ldf\relax

33 ⟨+greek⟩\input greek.ldf\relax
34 ⟨+irish⟩\input irish.ldf\relax
35 ⟨+italian⟩\input italian.ldf\relax
36 ⟨+lsorbian⟩\input lsorbian.ldf\relax
37 ⟨+magyar⟩\def\CurrentOption{magyar}
38 ⟨+hungarian⟩\def\CurrentOption{hungarian}
39 ⟨+magyar | hungarian⟩\input magyar.ldf\relax
40 ⟨+norsk⟩\def\CurrentOption{norsk}
41 ⟨+nynorsk⟩\def\CurrentOption{nynorsk}
42 ⟨+norsk | nynorsk⟩\input norsk.ldf\relax
43 ⟨+polish⟩\input polish.ldf\relax
44 ⟨+portuges⟩\def\CurrentOption{portuges}
45 ⟨+portuguese⟩\def\CurrentOption{portuguese}
46 ⟨+brazil⟩\def\CurrentOption{brazil}
47 ⟨+brazilian⟩\def\CurrentOption{brazilian}

48 ⟨+portuges | portuguese | brazil | brazilian⟩\input portuges.ldf\relax

49 ⟨+romanian⟩\input romanian.ldf\relax

50 ⟨+russian⟩\input russianb.ldf\relax
51 ⟨+scottish⟩\input scottish.ldf\relax
52 ⟨+slovak⟩\input slovak.ldf\relax
53 ⟨+slovene⟩\input slovene.ldf\relax
54 ⟨+spanish⟩\input spanish.ldf\relax
55 ⟨+swedish⟩\input swedish.ldf\relax
56 ⟨+turkish⟩\input turkish.ldf\relax
57 ⟨+usorbian⟩\input usorbian.ldf\relax
58 ⟨+welsh⟩\input welsh.ldf\relax

Change History

bbcompat-1.2

General: Added .sty files and def-

      inition of \CurrentOption for

language definition files that are loaded by more than one option. . . . . . . . . . . . . . . . . . . . . 1 bbcompat1.2

General: Added the check for

\LdfInit . . . . . . . . . . . . . . . . 1 bbxompat-1.2

General: Added welsh.sty . . . . . . 1

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\@ . . . . . . . . . . . . 3, 4, 6

\@undefined . . . . . . . . 1

A

\atcatcode . . . . . . 3, 6

C

\catcode . . . . . . . 3, 4, 6 \chardef . . . . . . . . . . 3

\CurrentOption . . . .

9–11, 19, 20, 25, 26, 29–31, 37, 38, 40, 41, 44–47

D \def . . . . . . . . . . . 2,

9–11, 19, 20, 25, 26, 29–31, 37, 38, 40, 41, 44–47

F \fi . . . . . . . . . . . . . . 8

I \ifx . . . . . . . . . . . . . . 1 \input 5, 12–18, 21–24,

27, 28, 32–36, 39, 42, 43, 48–58

L \LdfInit . . . . . . . 1, 2, 7 \let . . . . . . . . . . . . . . 6

R \relax . . . . . . . . . 4–

6, 12–18, 21–24, 27, 28, 32–36, 39, 42, 43, 48–58